OpenStudy (anonymous):

find the x-interceps and vertex of the function y=x^2+2x+6

OpenStudy (anonymous):

help please!!

OpenStudy (anonymous):

first factor the problem, that will give you the x intercepts.

OpenStudy (anonymous):

okay so (x+3)(x+2)? i think i cant figure out how to factor it correctly

OpenStudy (anonymous):

so you can't factor it correct?

OpenStudy (anonymous):

i factored it buut im not sure if im right

OpenStudy (anonymous):

It's not, that's because you can't factor it. The parabola does not cross the x axis

OpenStudy (anonymous):

oh! so it doesnt have any x-intercepts?

OpenStudy (anonymous):

right, or at least no real ones, but that's another lesson.

OpenStudy (anonymous):

okay so how do you find the vertex? because dont you have to factor it first?

OpenStudy (anonymous):

do you remember the quadratic formula?

OpenStudy (anonymous):

yeah. do i put the equation in the quadratic formula and solve it?

OpenStudy (anonymous):

not exactly, but plug in the numbers for \[\left(\begin{matrix}-b \\ 2a\end{matrix}\right)\]

OpenStudy (anonymous):

this will give you the x of the vertex, (x,y)

OpenStudy (anonymous):

okay so -1?

OpenStudy (anonymous):

right, so plug that back into the equation for the y value

OpenStudy (anonymous):

okay so the vertex is (-1,5)?

OpenStudy (anonymous):

right. so x=no solution and vertex=(-1,5)

OpenStudy (anonymous):

omg! thanks. can you help me with more? i suck at algebra?

OpenStudy (anonymous):

sure what other questions do you have?

OpenStudy (anonymous):

find the x-corrdnate of the vertex of the function below you can enter just a number. y=x^2+10+21

OpenStudy (anonymous):

well can you factor it?

OpenStudy (anonymous):

yeah. (x+7) (x+3)

OpenStudy (anonymous):

so you can either average the x intercepts or use the equation I showed you earlier.

OpenStudy (anonymous):

so the vertex is -5? so.....(-5,-4)?

OpenStudy (anonymous):

or wait....it = -5?

OpenStudy (anonymous):

sorry, I think I lost you for a second. Are you still here?

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

but I got the right answer for that question. now im on a different one.

OpenStudy (anonymous):

yep that's right, do you still need help?

OpenStudy (anonymous):

im not sure. are you leaving?

OpenStudy (anonymous):

no, I'm here if you need me.

OpenStudy (anonymous):

okay thanks so much!!!

OpenStudy (anonymous):

no problem. hey how would I find you again if I just quickly answered another question?

OpenStudy (anonymous):

um you can become my fan. and then i can message you when your online kind of like facebook

OpenStudy (anonymous):

Okay, officially a fan. ask me any question any time. Except for spelling.